# MSCCL++: Rethinking GPU Communication Abstractions for Cutting-edge AI Applications

### v1

@@ 4.2.3 SwitchChannel
A SwitchChannel provides primitives for performing collective operations among GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast. A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. We now discuss the implementation of these primitives for NVIDIA NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using the NVLink SHARP (NVLS) technology [23]. However, we believe that these implementations can be generalized to other hardware in the future.
@@ 4.3 DSL Implementation
MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level. It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back. MSCCL++ DSL is an extension of the MSCCLang [10] that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time, etc.
MSCCL++ DSL introduces a few limitations over the Primitive API. First, as the DSL introduces higher level instructions than the Primitive API, it may not be the most efficient way to implement. In our evaluation, DSL-written algorithms perform 3% worse than the hand-written ones on average, and is up to 18% worse in a corner case. Second, since a DSL-written algorithm can be run only by the provided executor kernel, it may not be straightforward to merge with other computation kernels if needed. Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
4.4 Collective API and Algorithms
MSCCL++ Collective API is the highest level of API that MSCCL++ supports, and it is the same as the original NCCL API. It is implemented as a wrapper of the MSCCL++ DSL API, which means that users can install a DSL-written algorithm to be called by the NCCL API. Users can also use the default algorithms already provided by the API. At runtime, for each communication, MSCCL++ automatically selects the best performing algorithm.
To be specific, we provide high-level description of the AllReduce algorithms we implement using the MSCCL++ DSL API below. Each algorithm can be implemented in multiple ways for optimization depending on hardware and message size. For example, we implement eight different versions of the two-phase all-pairs (2PA) algorithm explained in a following paragraph, and we omit the details for brevity.
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficient than baselines by removing unnecessary memory copies and synchronization as described in Section 2.2.2.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collect and reduce $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) [31]. Two-phase algorithms are more bandwidth-efficient and conducts less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple versions with combinations of options: MemoryChannel or PortChannel, LL or HB protocol, and put-based or get-based. Using MSCCL++, we can optimize 2PA in various ways that are not possible with baselines. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space, as explained in Section 2.2.2. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction comparing to baselines that read data from different GPUs one-by-one, which syncrhonizes each reduction step.
3. Two-phase Hierarchical (2PH). Hierarchical algorithms exchange minimal data across nodes and do local collectives in each node to complete the operation. It can be faster than all-pairs by reducing the data traffic crossing the nodes. 2PH is a two-phase algorithm that performs ReduceScatter and AllGather in a hierarchical manner each. We use it for multi-node collectives and implement two versions. The first version is for small messages using LL protocol. Each node conducts local ReduceScatter that splits the data into the number of GPUs in a node. This requires to send more data across nodes and introduce redundant reduction, but it is faster for small messages by reducing synchronization steps. Cross-node communication is done in all-pairs manner. To utilize inter-/intra-node links at the same time, we pipeline the local collective with cross-node communication to overlap. The second version of 2PH is for large messages using HB protocol. Unlike the first version, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs. Similar to the first version, it performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner.
4. Two-Phase All-Pairs Multimem (2PAM). 2PAM is a specialized algorithm for NVIDIA H100 GPUs that leverages the in-network, switch-based aggregation and multicast capabilities by the NVLink SHARP (NVLS) technology [23]. This algorithm simply walks through all data elements one-by-one in a loop to call two NVLS multimem instructions, one for in-network cumulating one element from all GPUs, and another for broadcasting the cumulated element.
4.5 Cross-hardware Implementation
This section introduces our experience in developing MSCCL++ across different hardware platforms and how we reduce development efforts for future hardware.
4.5.1 Quick Support for New GPUs
Extending MSCCL++ to support new hardware features is much easier compared to doing so in existing communication libraries, because the primitive interface is directly exposed to the user. This design accelerates the overall development because (1) the library developers only need to add a shallow layer of abstraction over low-level code that features key functionalities, and (2) the algorithm developers can write clever optimizations and fine-tuning with precise control of the hardware efficiency.
We provide an example of supporting the NVIDIA H100 GPU in MSCCL++, which delivers the hardware-based collective operations using NVSwitch (explained in Section 4.2.3). Since this feature introduced a very different concept from previous communication interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than 2.2$\times$ in average for message sizes of 1KB - 1GB.
As another example, we have developed MSCCL++ to support the AMD MI300x GPU, while the previous version supported only NVIDIA GPUs. The development took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB. It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL. This is possible because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API. This significantly reduces the effort to maintain and develop MSCCL++.
4.5.2 Managing Memory Consistency
Besides the proposed abstractions, MSCCL++ also puts particular considerations on its low-level implementation to reduce dependencies on certain hardware, thus reducing technical debt. As an example, we introduce our approach toward managing memory consistency. Ensuring memory consistency is critical as a communication library. In particular, fast-evolving chips and their compilers often consist of various bugs or undocumented behaviors that can lead to breaking memory consistency, which is often hard to detect and fix. This is a solid problem that prevents the quick adoption of the latest hardware techniques.
From our production efforts to tackle this problem, we found that the most effective and efficient way to ensure memory consistency is relying on the atomic operations, rather than leveraging other consistency models provided by hardware or compiler features. This makes it easier to blame the hardware or compiler when we find unexpected behaviors, [^8] Indeed, we found a bug of an atomic instruction in CUDA 12.x NVCC, which was reported to and fixed by NVIDIA. and also makes MSCCL++ more portable and maintainable. For example, we use atomic instructions that comply with the C++11 memory model instead of volatile instructions. As another example, for signal over RDMA, one may simply send a separate RDMA write request to send a flag to the target GPU, assuming that the underlying hardware and stacks are configured and working correctly to ensure ordered write. However, instead, MSCCL++ uses the atomic fetch-and-add operation of RDMA to ensure consistency from the software perspective (the impact on performance is negligible small).
@@ 4.5.2 Managing Memory Consistency
Table 1: List of environments used for evaluation.
5 Evaluation
@@
Environments. All presented numbers are collected from one of the environments listed in Table 1. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)), and inter-node InfiniBand links (one NIC per GPU). All NICs are connected to a single InfiniBand networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
@@
Baselines. We compare MSCCL++ with existing collective communication libraries, including NCCL 2.23.4 [21], RCCL 2.20.5 [4], [^9] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL. and MSCCL 2.23 [19]. MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms. Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
5.1 Collective Communication
@@
We present collective communication performance of MSCCL++ and compare it with that of the baselines. We implement various GPU kernels based on the algorithms in Section 4.4, and we present the best number among all implementations for each message size as we do for NCCL, RCCL, and MSCCL as well. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS [23]), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each buffer size [20]. We leverage NCCL’s user buffer registration (i.e., ncclMemAlloc[26] and the CUDA/HIP Graph features for best performance.
@@
For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^10] Defined as the message size divided by the latency.) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode [30]), while the large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing, a.k.a. prefill [30]).
Figure 8: AllReduce, A100-40G, single-node (8 GPUs).
Figure 9: AllReduce, A100-40G, 2-node (16 GPUs).
Figure 10: AllReduce, A100-40G, 4-node (32 GPUs).
Figure 11: AllReduce, H100, single-node (8 GPUs).
Figure 12: AllReduce, MI300x, single-node (8 GPUs).
We compare the AllReduce performance over A100-40G nodes in Figure 8, 9, and 10. All data points are validated multiple times with separate trials. MSCCL++ outperforms baselines for both small messages (up to 3.5x and 2.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.6x and 1.4x faster over NCCL and MSCCL, respectively). MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms. The gain is especially big for small message sizes where MSCCL uses all-pairs algorithms, while NCCL uses the ring algorithm that is worse in terms of latency. The gap between MSCCL and NCCL shrinks for very large message sizes where the performance is limited by the link bandwidth. MSCCL++ provides further benefits over MSCCL by implementing the algorithms in a more efficient way via the primitive API. All MSCCL and MSCCL++ numbers in the figures use the same collective algorithms and differ only by implementation in all cases. Even if MSCCL and MSCCL++ separately select the best-performing algorithm for each case, they come up with the same algorithm. Therefore, the gap between MSCCL and MSCCL++ directly shows the benefits of the proposed primitive API.
@@
Figure 11 and 12 compare the single-node AllReduce performance over H100 and MI300x nodes. Similarly, MSCCL++ outperforms baselines by up to 3.8x and 2.2x for small and large message ranges, respectively. This shows that MSCCL++ is effective across different GPU architectures and links. The speedup over NCCL/MSCCL for large message sizes on H100 is especially interesting, because the algorithm used by MSCCL++ here is 2PAM (Section 4.4), which simply calls two NVLS multimem instructions element-wise (one for reduction and one for broadcast) in a loop. This shows that NCCL primitives are built with a big overhead of unnecessary components that can be avoided in MSCCL++.
5.2 LLM Inference Acceleration
Figure 13: Speedup for decodes with tensor parallelism = 8
@@ 5.2 LLM Inference Acceleration
6 Related Work

### v2

@@ 4.2.3. SwitchChannel
A SwitchChannel provides primitives for performing collective operations among GPUs. These operations usually require specialized hardware support. For example, NVIDIA NVLink 4.0 connects all H100 GPUs to a single NVSwitch and this NVSwitch can perform collective operations including reduce and broadcast. A SwitchChannel in MSCCL++ provides two primitives: reduce to add the corresponding elements of buffers residing on different GPUs and broadcast to send elements from a buffer on a GPU to all other GPUs. We now discuss the implementation of these primitives for NVSwitch, which provides in-network, switch-based aggregation and multicast capabilities using NVLink SHARP (NVLS) technology (NVIDIA, 2025c). However, we believe that these implementations can be generalized to other hardware in the future.
@@ 4.3. MSCCL++ DSL API and Executor
MSCCL++ DSL API is a Python-based language that allows describing a communication algorithm at a high level. It converts the algorithm description into a sequence of instructions that can be executed by the DSL Executor, which is a GPU kernel that reads and runs instructions back-to-back. MSCCL++ DSL is an extension of the MSCCLang (Cowan et al., 2023) that supports new instructions based on the MSCCL++ Primitive API and lifts a few restrictions of the language; e.g., we enable a single GPU thread block to access multiple GPUs at the same time.
4.4. MSCCL++ Collective API and Kernels
MSCCL++ Collective API is the simplest API that MSCCL++ offers, and it is the same as the popular NCCL API. We implement this API using kernels written in the MSCCL++ Primitive API. Users can either use these default kernels or plug-in their own algorithms written using the MSCCL++ DSL or Primitive APIs. We now briefly describe a few AllReduce algorithms included in the default kernels to: (i) provide ideas on low-level considerations for implementing high-performance GPU communication kernels, and (ii) to show how low-level interfaces can be useful for customization on different workloads.
@@ 4.4. MSCCL++ Collective API and Kernels
Table 1. List of environments used for evaluation.
1. One-phase All-pairs (1PA). In all-pairs algorithms, all GPUs concurrently broadcast their own local data to all other GPUs. By one-phase, the algorithm sends all local data to all other GPUs, so that the reduction is done in a single phase. It is suitable for small message sizes where the synchronization overhead between GPUs is more critical than the redundant reduction and data traffic. As we use the 1PA algorithm only for very small message collectives within a single node, we implement only a single version of 1PA that uses the LL protocol with MemoryChannel. MSCCL++ can implement it much more efficiently than other libraries by removing synchronizations.
2. Two-phase All-pairs (2PA). By two-phase, the algorithm splits the AllReduce into two phases: the first for ReduceScatter (each of $N$ GPUs collects and reduces $1/N$ of the data) and the second for AllGather (each GPU broadcasts the reduced data to all other GPUs) (Rabenseifner, 2004). Two-phase algorithms are more bandwidth-efficient and conduct less reduction than one-phase algorithms. In 2PA, the ReduceScatter and AllGather phases are done in the all-pairs manner each. We use 2PA for single-node collectives and implement multiple variants using PortChannel, MemoryChannel with LL or HB protocol, or SwitchChannel. Using MSCCL++, we can optimize 2PA in various ways that are not possible in existing libraries. For example, for up to a few MBs of messages, we exploit rotating buffers to reduce synchronization at the cost of using more memory space. As another example, we can let a single thread group read data from multiple other GPUs at the same time. This allows efficient data reduction compared with other libraries that read data from different GPUs one-by-one, which synchronizes each reduction step.
3. Two-phase Hierarchical (2PH). Hierarchical algorithms exchange minimal data across nodes and do local collectives in each node to complete the operation. It can be faster than all-pairs by reducing the data traffic crossing the nodes. 2PH is a two-phase algorithm that performs ReduceScatter and AllGather in a hierarchical manner each. We use it for multi-node collectives and implement two variants. The first version is for small messages using the LL protocol. Each node conducts local ReduceScatter that splits the data into the number of GPUs in a node. This requires sending more data across nodes and introduces redundant reduction, but it is faster for small messages by reducing synchronization steps. Cross-node communication is done in all-pairs manner. To utilize inter-/intra-node links at the same time, we pipeline the local collective with cross-node communication to overlap. The second version of 2PH is for large messages using HB protocol. This variant also performs all-pairs cross-node communication and local ReduceScatter in a pipelined manner. Unlike the first variant, to utilize the link bandwidth efficiently, the number of data chunks is the same as the number of GPUs.
4.5. Portability and Implementation Effort
MSCCL++ is designed to be portable. Code using MSCCL++ collectives, algorithms using the MSCCL++ DSL, and kernels using the MSCCL++ primitives are portable as long as the underlying MSCCL++ Primitive API supports that hardware. Our design makes it easy to add support for new hardware. As an example, MSCCL++ initially supported only NVIDIA GPUs; to add support for AMD MI300x GPUs, our implementation took only 7 weeks for one developer: 3 weeks for basic AMD GPU support and 4 weeks to develop new AllReduce algorithms that outperform RCCL/MSCCL for message sizes of 1KB - 1GB. It is especially remarkable that the AMD-specific code in MSCCL++ (excluding makefiles and algorithms) is only less than 10 lines of code, while RCCL, the hard-forked NCCL for AMD GPUs, is substantially diverged from NCCL. MSCCL++ could achieve this because the low-level API of AMD GPUs (i.e., HIP) is almost the same as that of NVIDIA GPUs (i.e., CUDA), and the MSCCL++ Primitive API is only a shallow abstraction on top of the low-level API. This significantly reduces the effort to maintain and develop MSCCL++.
The MSCCL++ design makes it easy to extend the Primitive API to support new hardware features. For example, NVIDIA NVLink 4.0 introduced support for multimem instructions for switch-based aggregation and multicast. Since this feature introduced a very different concept from previous interfaces (i.e., PortChannel or MemoryChannel), we developed the SwitchChannel interface to support this feature. The development took only 8 weeks for two developers, including learning the basic usage of this feature, abstracting the feature as a new type of channel (i.e., SwitchChannel), and finally developing a new AllReduce algorithm using SwitchChannel that outperforms NCCL/MSCCL by more than $\times2$ in average for message sizes of 1KB - 1GB.
5. Evaluation
@@
Environments. All presented numbers are collected from one of the environments listed in Table 1. Each node is equipped with 8 GPUs per node (either A100, H100, or MI300X), intra-node links between GPUs (either NVLink or Infinity Fabric (a.k.a. xGMI)), and inter-node IB links (one NIC per GPU). All NICs are connected to a single IB networking switch. NVIDIA GPUs use CUDA 12.4, while AMD GPUs use ROCm 6.2. For brevity, the rest of this paper refers to each environment by the name of GPU.
@@
Baselines. We compare MSCCL++ with the state-of-the-art collective communication libraries, including NCCL 2.26.2 (NVIDIA, 2025a), RCCL 2.20.5 (AMD, 2025), [^5] As RCCL partly adopts MSCCL, to prevent confusion, we refer to MSCCL-disabled RCCL. and MSCCL 2.23 (Microsoft, 2025a). MSCCL borrows the stack implementation of NCCL/RCCL and supports custom algorithms. Therefore, MSCCL performs exactly the same as NCCL/RCCL if they use the same algorithm.
5.1. Collective Communication
@@
To compare the performance of collective communication, we implement the AllReduce algorithms described in Section 4.4 using the MSCCL++ DSL API. We also implement various AllGather algorithms in a similar manner, and we omit the implementation details. We present the best number among all implementations for each message size and environment. All NCCL, RCCL, and MSCCL numbers are fine-tuned for each environment and message size by adjusting their environment variables, such as the number of channels (affects the number of threads), chunk size (affects the size of data to be transferred at once), type of algorithm (such as ring, tree, or NVLS (NVIDIA, 2025c)), the topology (XML file that describes the intra-node link topology of GPUs), etc. For MSCCL, we use the fastest algorithm for each message size (Microsoft, 2025b). We use NCCL’s user buffer registration API (i.e., ncclMemAlloc(NVIDIA, 2025f) and the CUDA/HIP Graph APIs for best performance.
@@
For visibility, we separate the range of message sizes into small (up to 1MB, presented latency) and large (1MB and above, presented algorithm bandwidth (AlgoBW) [^6] Defined as the message size divided by the latency.) in the figures. The small message sizes represent inference scenarios (such as LLM token sampling a.k.a. decode (Pope et al., 2023)), while large message sizes represent both training and inference scenarios (such as gradient accumulation during back-propagation, or LLM prompt processing a.k.a. prefill (Pope et al., 2023)).
Results. Figure 8 compares the AllReduce performance over A100-40G. MSCCL++ outperforms baselines for both small messages (up to 4.2x and 3.1x faster over NCCL and MSCCL, respectively) and large messages (up to 1.8x faster over both NCCL and MSCCL). Similarly, Figure 9 compares the AllGather performance over A100-40G. For small messages, MSCCL++ outperforms NCCL and MSCCL by up to 5.4x and 2.3x, respectively, and for large messages, MSCCL++ outperforms NCCL and MSCCL by up to 1.8x and 1.4x, respectively. For AllGather, we observe a few cases where MSCCL slightly outperforms MSCCL++ by up to 8.0%, which is introduced by a performance bug of the DSL API, not by our algorithm implementation. For AllReduce, MSCCL++ is always the fastest.
Figure 8. AllReduce, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
Gain Breakdown. The performance differences can come from both high-level differences in algorithms (i.e., design of data flow routes and pipelines) and low-level optimizations (i.e., how to transfer data, synchronize, and manage buffers). The former benefit is already delivered by MSCCL that enables the customization of algorithms over the NCCL communication stack. Therefore, the key benefits of MSCCL++ lie in the latter by exposing the primitive functionalities to users. Since MSCCL and MSCCL++ use the same high-level algorithm for all cases in Figure 8 and 9, the performance gap between the two comes purely from the benefits of MSCCL++.
NCCL vs MSCCL. MSCCL is faster than NCCL in most cases, and all the benefit comes from better algorithms customized for our environments. The gain is especially big for small message sizes, where MSCCL uses an all-pairs algorithm (which is not supported by NCCL), while NCCL uses the ring algorithm that is worse in terms of latency. For very large message sizes $\geq$ 256MB in multi-node cases, MSCCL achieves higher bandwidth than NCCL because it uses a hierarchical algorithm (2PH), which is more bandwidth-efficient.
Figure 9. AllGather, A100-40G. MnNg means M nodes and N GPUs. X-axis is message size (at the bottom of the figure).
@@
DSL vs Primitive. From our 9-month experience using the MSCCL++ DSL API, we find that it reduces development time from weeks to days, compared to using the Primitive API directly. The DSL helps automate initialization, check for mistakes, and provide performance analysis. We have implemented both the DSL and Primitive versions up to 2-node experiments to verify that their performance is similar, while 4-node experiments are done only using the DSL. However, as the DSL introduces higher level instructions than the Primitive API, DSL versions perform 3% worse than the Primitive versions on average, and are up to 18% worse in one corner case. Despite limitations, MSCCL++ DSL is still useful for quick prototyping of collective communication and easy-to-understand algorithm description.
Collective vs Primitive. For AI inference workloads, we implement the best algorithms in our collective kernels using the SpeCL Primitive API, so there is no performance difference.
Figure 10. Speedup for decodes with tensor parallelism = 8
5.2. LLM Inference
@@
5.3. Cross-Hardware Support
To show the effectiveness of MSCCL++ on different hardware platforms, Figure 11 and 12 compare the single-node AllReduce performance on H100 and MI300x.
[…]
Figure 11. AllReduce, H100, single-node (8 GPUs).
Figure 12. AllReduce, MI300x, single-node (8 GPUs).
MI300x. MSCCL++ outperforms RCCL and MSCCL by up to 3.8x and 1.9x for small messages, and by up to 2.2x and 1.6x for large messages, respectively. Our algorithm implementations take special considerations to best utilize the Infinity Fabric that peer-to-peer connects all GPUs in a node, unlike NVLink that connects all GPUs to a switch. This means, for best link utilization, we need to copy data to all peers at the same time as much as possible, unlike for NVIDIA GPUs where we can copy data to each peer one-by-one back-to-back. Using the MSCCL++ DSL, this is as easy as changing the order of two nested for loops, one that walks through data elements and the other that walks through different peers.
6. Related Work